OpenStudy (anonymous):

What are the solutions of 6x2 − x − 15 = 0? please i really dont understand

OpenStudy (anonymous):

Do you know how to factor by grouping?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

ax^2+bx+c=0 6x^2-x-15=0 numbers that add together to get B and multiply to get AC AC: 6*-15=90 B: -1 What two numbers will add to get you B and multiply to get you AC?

OpenStudy (anonymous):

b=-2+1?

OpenStudy (anonymous):

No, the two numbers that will multiply to get you -90 and add to get you -1.

jhonyy9 (jhonyy9):

Ruka there AC =6*(-15)= -90 and not 90 ATTENTION please !!!

OpenStudy (anonymous):

@jhonyy9 ATTENTION please !!! READ THE REPLY ABOVE YOU BEFORE YOU COMMENT.

jhonyy9 (jhonyy9):

ok - sorry !!!